A peace of research equipment is expected to require an investment of 18000$ with 6000$ committed now and next year 7000$ while the remaining 6K$ will be paid out at the end of year 2. Annual operating costs for the system are expected to start in the first year and continue at 800$ per year. The life of the equipment is 8 years with a salvage value of 500$. Calculate the CR and AW. Values for the system, if the corporate MARR is 14.i ii% per year
1
Expert's answer
2021-08-29T16:52:58-0400

A piece of research equipment is expected to require an investment of $18,000, with $6,000 committed now and next year $7000 while the remaining $6K will be pay out at the end of year 2. Annual operating costs for the system are expected to start in the first year and continue at $800 per year. The life of the equipment is 8 years with a salvage value of $500. Calculate the CR and AW values for the system, if the corporate MARR is 14.32% per year.



Year wise cash flow are given

Present value "= \frac{cf }{ ( 1+i)^n}"

For year 0 "= \frac{-6000}{ (1+14.32 \%)^0} = -6000"

For year 1 "= \frac{-6000}{ (1+14.32 \%)^1}= -5248. 42"

For year 2 "= \frac{-6000}{(1+14.32 \%)^2}= -4590. 99"

For year 8 "= \frac{-500}{ (1+14.32 \%)^8}= 171.39"

Capital recovery "= -6000-5248.42-4590.99+171.39"

Capital recovery "= - 15668.02"

Annual worth "= -15668.02- 800 = -16468.02"